OpenStudy (kmullis6):

Am I right? Solve for x: 3(x + 1) = -2(x - 1) + 6. 1 4 5 25

OpenStudy (kmullis6):

I think it's 1.

OpenStudy (kmullis6):

I wanna make sure I'm doing this right. /:

OpenStudy (johnweldon1993):

You are correct :)

OpenStudy (kmullis6):

@johnweldon1993 ok thanks, can you help me on one more?

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (kmullis6):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432

OpenStudy (johnweldon1993):

Alright so that equation gives us how many were between 30 and 40...so we jut need to solve that for 'x' x = 0.8x + 72 Combine like terms...subtract 0.8x from both sides of the equation x - 0.8x = 72 .2x = 72 x = (72/.2)

OpenStudy (kmullis6):

Sorry, I'm terrible at this lol but where did the .2x come from?

OpenStudy (johnweldon1993):

No dont worry about it :) so we had x - 0.8x = 72 Well 1x - 0.8x = (1 - 0.8)x 1 - 0.8 = 0.2 right?

OpenStudy (kmullis6):

ohhh ok my bad

OpenStudy (kmullis6):

Ok I get it now thanks so much (: @johnweldon1993

OpenStudy (johnweldon1993):

Of course :)